A student studying for a vocabulary test knows the meanings of 16 words from a list of 20 words. If the test contains 10 words from the study list, what is the probability that at least 8 of the words on the test are words that the student knows? (Round your answer to three decimal places.)
Answer #1

Number of ways = 20C10

Number of ways for 8 words  = 16C8 * 4C2

Number of ways for 9 words  = 16C9 * 4C1

Number of ways for 10 words  = 16C10 * 4C0

P(X>=8) = P(X=8) +P(X=9) +P(X=10)

=16C8 * 4C2 / 20C10 + 16C9 * 4C1 / 20C10 + 16C10 * 4C0 / 20C10

= 0.417956656 + 0.247678019 + 0.043343653

=0.709

Please revert in case of any doubt.

Please upvote. Thanks in advance